P1-2. A movie company is making a new film. For the main role, they hired Johhny and he will receive a salary of $20,000,000 + 5% from admissions. The producer normally receives 40% of total paid admission whenever the movie is shown. Cost of production i

A movie company is making a new film. For the main role, they hired Johhny and he will receive a salary of $20,000,000 + 5% from admissions. The producer normally receives 40% of total paid admission whenever the movie is shown. Cost of production is $45,000,000. Also, a study of past films indicates that the producer can expect revenues from TV rights to be 12.5% of the producer's revenues from admissions.


Questions:

a) Considering the additional information about the sale of television rights, at what level of receipts to the producer will Blockbusters break even under Johny's contract terms ($20,000,000 plus 5% of the producer’s receipts)?


b) Johny asked for a special contract with a salary of $10,000,000, plus 15% of the receipts to the producer from admissions, plus 10% of the revenues from TV rights. At what level of receipts to the producer would Johny earn the same total income under both contracts?
Answer #1

Let's get the facts first

1) Producer will earn revenue from admission and from TV rights. Revenue from TV rights = 1/8th of revenue from admission

2) His cost of production excluding payment to Liam = $ 45,000,000

3) Liam has offered 2 proposals - Normal (N) and Special (S)

4) Under N, producer will pay him a fixed $ 20,000,000 plus 5% of revenue from admission

5) Under S, producer will pay him a fixed $ 10,000,000 plus 15% of revenue from admission plus 10% of revenue from TV rights

6) We have to find BEP to producer under both alternatives. BEP means total revenue = Total cost

7) If we assume revenue from admission to be 'x'. Revenue from TV = x/8. Total revenue = x+x/8 = 9x/8

8) Payment to Liam under N = 20,000,000 + 5x/100

9) Total cost under N = 45,000,000 + 20,000,000 + 5x/100

10) Payment to Liam under S = 10,000,000 + 15x/100 + 10/100×(x/8) = 10,000,000 + 15x/100 + x/80

11) Total cost under S = 45,000,000 + 10,000,000 + 15x/100 + x/80

12) BEP when revenue = cost under each alternative.

Solved in 3 sheets uploaded

BEP under N = Revenue from admission = $ 60,465,116

BEP Under S = Revenue from admission = $ 57,142,857

BEP is lower under S. Hence S is less risky to producer. Hence producer should prefer Special contract

Payment to Liam is higher under N ($ 23,023,256) as against S ( $ 19,285,714). Hence Liam should prefer Normal contract

1/3 BEP . Normal Contract Revenne Revenne from admission = Revenue from Tu rights = Total revenues (x+ Total cost = 45,00 00

Ele BEP - Special Revenue from admission TV rights Total revenue xxx 층 Pay to Liam= 10.000.000 + 15 رع = 92 que Tooxxt to TOOverification as Revenue from admission Revenne from TV rights t Y8th) 57.142857 7142857 64.285,714 Total revenue pay. to Liam